# Frustrated Quantum Magnetism on Complex Networks: What Sets the Total Spin

### v2

@@ Abstract
Figure 6: (a) Wheel graph with $2$ central spins and $N_{b}=6$ outer spins. (b) Phase diagram for $N_{b}=30$ and rescaled couplings $\tilde{J}_{b}\coloneqq 4J_{b}/(JN_{c})$ and $\tilde{J}_{c}\coloneqq J_{c}N_{c}/(JN_{b})$. Green arrow shows a discontinuous transition from a state where central and outer spins are oppositely aligned to a state where both add up to zero. Blue arrow shows a continuous transition where the net outer spin $S_{b}$ falls in steps of $1$ at a rate $\sim N_{b}(\tilde{J}_{b}-1)^{-1/2}$ (inset). This divergence is absent for classical Heisenberg spins (dashed curve), for which $S_{\text{total}}=0$ for $\tilde{J}_{b}\tilde{J}_{c}>1$ (solid curve).
@@
To understand its behavior, note that the spokes favor both the outer spin $S_{b}$ and the central spin $S_{c}$ to be large but anti-aligned, i.e., $S_{\text{total}}=|S_{b}-S_{c}|$, whereas $J_{b}$ and $J_{c}$ want to suppress $S_{b}$ and $S_{c}$, respectively. For simplicity, let us consider $N_{c}=2$ and $N_{b}$ even. Then the central spins form either a singlet or a triplet, depending on how strongly the central bonds compete with the spokes. For $J_{c}N_{c}^{2}\gg JN_{b}N_{c}$ we expect $S_{c}=0$, which reduces the network to just the ring, yielding $S_{b}=0$ and $S_{\text{total}}=0$. In contrast, for $J_{c}N_{c}^{2}\ll JN_{b}N_{c}$ we expect $S_{c}=1$, which leads to a competition between the spokes and the outer bonds: For $J_{b}N_{b}\ll JN_{b}N_{c}$ the outer spins align, so that $S_{b}=N_{b}/2$ and $S_{\text{total}}=N_{b}/2-1$, whereas for $J_{b}N_{b}\gg JN_{b}N_{c}$ they cancel, giving $S_{b}=0$ and $S_{\text{total}}=1$. Thus, by varying $J_{b}$ and $J_{c}$ one should be able to stabilize any total spin between 00 and $N_{b}/2-1$.
These intuitions are confirmed by solving the system exactly using the Bethe Ansatz solution for the Heisenberg ring [65, 66] (see Supplement [49] for details). For $N_{c}=2$ this gives the phase diagram in Fig. 6(b). As expected, we find an extended antiferromagnetic region where $S_{b}=S_{c}=0$ and an extended ferromagnetic region where both spins are maximum but oppositely aligned. The two regions share a border where the magnetization collapses abruptly (green arrow). On the other hand, for small $J_{c}$ the outer spin falls in steps of $1$ with $J_{b}$, exhibiting a power-law diverging susceptibility at $J_{b}=JN_{c}/4$. Interestingly, this blow up is absent for classical Heisenberg spins (see Supplement [49]), for which $S_{\text{total}}$ is generally higher. For odd $N_{c}$ the center cannot form a singlet, so there is always a competition between $J$ and $J_{b}$, leading to a stepwise decrease of $S_{\text{total}}$ (see Supplement [49] for the phase diagram for $N_{c}=3$).
Outlook.—Perhaps the most promising platform for realizing our setup is ion traps, for which there are concrete protocols for engineering arbitrary pairwise Heisenberg coupling with existing technology for several tens of qubits [25, 26, 27]. A large class of networks can also be fabricated in photonic platforms [29, 30], where Heisenberg interactions may be simulated digitally [28, 35] or using tailored light-matter coupling [67]. Our results support the general view that hubs promote cohesion across a network [4] (e.g., ferromagnetic ordering [8, 68, 69, 70]), while pointing out the importance of assortativity [6, 71].
At the same time, they raise a number of open questions: (1) Why is $S_{\text{total}}$ insensitive to frustration? This is far from obvious and likely depends on the mechanism for magnetic interactions. In fact, preliminary calculations show that frustration strongly affects $S_{\text{total}}$ when magnetism arises from the delocalization of a doped carrier [72, 73]. (2) A related and perhaps the most crucial question concerns the nature of the ground state. Studies of classical Ising spins have found a glassy phase on small-world and scale-free networks [8, 74, 75, 76]. Are quantum fluctuations in the Heisenberg model strong enough to stabilize a spin liquid in the absence of a lattice symmetry [77]? (3) What are the dynamical and ergodicity properties and how are these affected by the network topology? (4) Is the magnetic order sensitive to motifs [78] other than hubs, or to community structures that regulate cooperative phenomena such as synchronization [71]? The latter arises in cluster magnetism [79]. (5) How does the physics change with bond disorder [8] or anisotropy? (6) As natural interactions decay with distance, it is useful to model spatial networks [9], which may be designed with Rydberg atoms in tweezers [80, 81]. Finally, we hope our findings herald broader explorations of quantum many-body networks beyond frustrated magnetism.
We thank Sitabhra Sinha and Sthitadhi Roy for useful discussions and comments. SD acknowledges use of the HPC facility at MPIPKS, Dresden.
@@
Random trees
A tree has a bipartite structure [Fig. S1(a)], for which $S_{\text{total}}=|N_{A}-N_{B}|/2$, where $N_{A}$ and $N_{B}$ are the numbers of spins in the two sublattices. Figure S1(b) shows that both the average and the spread of $S_{\text{total}}$ over random trees with $N$ sites and $N-1$ bonds scale as $\sqrt{N}$.
Figure S1: (a) Tree with $N=30$ sites. Horizontal lines show the two sublattices. (b) The average and the standard deviation of $S_{\text{total}}$ over $\mathcal{N}=1000$ random trees as a function of $N$. Dashed lines show the fits $0.23\sqrt{N}$ and $0.17\sqrt{N}$, respectively.
Random regular graphs
In the main text we showed that a random regular graph, where each node has $k=3$ neighbors, almost always has $S_{\text{total}}=0$. Figure S2 shows the same holds for $k=4$ and $k=5$, further demonstrating that heterogeneity is essential for nonzero total spin.
@@
Small-world networks
One can interpolate between a regular lattice and random graphs by randomly rewiring a small fraction of the bonds of the lattice; These “shortcuts” dramatically shorten the average path length between two nodes while maintaining a high clustering, resulting in a small-world character [1]. Figure S3 shows that this procedure leads to a slow increase of the average total spin, which is accompanied by a similar rise in the heterogeneity. Thus, we find no strong dependence of $S_{\text{total}}$ on the path length.
Figure S3: (a) Path length, (b) total spin and heterogeneity as a function of the rewiring probability $p_{r}$ for Watts-Strogatz graphs with $N=30$ and $\bar{k}=4$.
(Dis)assortative scale-free networks
In the main text we explained how tuning the assortativity of random graphs strongly affects $S_{\text{total}}$, which falls sharply with the assortativity coefficient $A$ [2]. Figure S4 shows the same is true for Barabási-Albert graphs with scale-free degree distribution ($p_{k}\sim 1/k^{3}$).
Figure S4: (a) Total-spin distribution of the most assortative ($\bar{A}=+0.17$) and the most disassortative ($\bar{A}=-0.78$) scale-free graphs with $N=30$ and $\bar{k}=3.8$. (b) Variation of $\bar{S}_{\text{total}}$ with the average assortativity. Dashed curve is for Ising spins.
Exact solution for the wheel
@@
In order to lower energy $\mathbf{S}_{b}$ and $\mathbf{S}_{c}$ point in opposite directions. Without loss of generality we take $\mathbf{S}_{b}$ along $+z$ and $\mathbf{S}_{c}$ along $-z$. The maximum magnitude of $\mathbf{S}_{c}$ is $N_{c}S$, where $S$ is the magnitude of each spin. To compare with qubits we set $S=1/2$ and write $\mathbf{S}_{c}=-(N_{c}/2)u\hat{\mathbf{z}}$, where $u\in[0,1]$. This central spin acts as an effective magnetic field for the Heisenberg ring [see Eq. (S1)], whose thermodynamic properties can be solved exactly [3]. The field wants to polarize the spins along $+z$, whereas the bonds prefer a Néel-type antiferromagnet. In the ground state, the spins orient along the polar and azimuthal angles
@@
where $\theta$ goes from 00 to $\pi/2$ as the field decreases, $\phi_{0}$ is an arbitrary constant, and $\alpha=\big{[}1-(-1)^{N_{b}}\big{]}\pi/(2N_{b})$. Substituting these into the Hamiltonian gives the energy
@@
$u=v=0\quad$ if $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}>1\;,$ (S4a)
$u=v=1\quad$ if $\text{if }\tilde{J}_{c}<1\text{ and }\tilde{J}_{b}^{\prime}<1\;,$ (S4b)
$u=1,v=1/\tilde{J}_{b}^{\prime}\quad$ if $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}<1\text{ and }\tilde{J}_{b}^{\prime}>1\;,$ (S4c)
$v=1,u=1/\tilde{J}_{c}\quad$ if $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}<1\text{ and }\tilde{J}_{c}>1\;,$ (S4d)
@@
Figures S5(b) and S5(c) show that the resulting phase diagram differs qualitatively for $N_{c}=2$ and $N_{c}=3$. As explained in the main text, this is because for odd $N_{c}$ one cannot have $S_{c}=0$, which leads to a stepwise fall of $S_{b}$ as $J_{b}$ is increased relative to $J$. To understand this feature more precisely we consider the energy as a function of $S_{b}$ for a given $S_{c}<S_{b}$,
@@
which is seen in Fig. S5(b) for $S_{c}=1$ and in Fig. S5(c) for $S_{c}=3/2,1/2$. No such divergence is present for classical spins [Fig. S5(a)].
Quantum vs. classical $S_{\text{total}}$
Figure S6: (Top panel) Correlation and (bottom panel) root-mean-square deviation between the net magnetizations of classical and quantum spins for random graphs (left column) and graphs generated by the growth model introduced in the main text (right column) with $N=30$ sites.

### v3

@@ Abstract
Figure 6: (a) Wheel with $N_{c}=2$ central spins and $N_{b}=6$ outer spins. (b) Phase diagram for $N_{b}=30$ where $\tilde{J}_{b}\coloneqq 4J_{b}/(JN_{c})$ and $\tilde{J}_{c}\coloneqq J_{c}N_{c}/(JN_{b})$. Green arrow shows a discontinuous transition from a state where central and outer spins are antialigned to a state where both sum to zero. Blue arrow shows a continuous transition where the net outer spin $S_{b}$ falls in steps of 11 at a rate $\sim N_{b}(\tilde{J}_{b}-1)^{-1/2}$ (inset). This divergence is absent for Ising spins (dotted) and classical Heisenberg spins (dashed). For the latter, $S_{\text{total}}=0$ for $\tilde{J}_{b}\tilde{J}_{c}>1$ (solid curve).
These intuitions are confirmed by solving the system exactly using the Bethe Ansatz for a Heisenberg ring [71, 72]. As shown in Fig. 6(b), the phase diagram features both abrupt and continuous transitions. Most notably, one finds a line of quantum critical points with power-law diverging susceptibility [70, 71, 72] at $J_{b}=JN_{c}/4$, which is absent for classical Heisenberg spins. (See SM [52] for derivations and more general comparisons between classical and quantum Heisenberg networks.) Note that Ising spins show only abrupt transitions from $S_{\rm total}=0$ to $S_{\rm total}\geq(N_{b}-N_{c})/2$ [Fig. 6(b) inset].
Outlook—Perhaps the most promising platform for realizing our setup is ion traps, for which there are concrete protocols for engineering arbitrary pairwise Heisenberg coupling with existing technology for several tens of qubits [28, 29, 30]. A large class of networks can also be fabricated in photonic platforms [32, 33], where Heisenberg interactions may be simulated digitally [31, 38] or using tailored light-matter coupling [73].
Our findings have parallels in classical network science, where heterogeneity has been found to promote cooperative phenomena. For instance, it greatly increases the critical temperature for ferromagnetic Potts models and reduces the synchronization threshold for Kuramoto oscillators [4, 13]. These correlate with a large ground-state total spin, which provides a natural way to synchronize by adding damping and a magnetic field that rotates the spins. Furthermore, we show that the impact of heterogeneity can be hugely enhanced by adding disassortative correlations (Fig. 4), and that quantum effects can yield a qualitatively different response (Fig. 6). These results strongly encourage the exploration of collective dynamical phenomena on quantum many-body networks.
They also raise a number of open questions on frustrated magnetism: (1) Why is the total spin generally insensitive to frustration level when all bonds are equal? This is far from obvious and does not hold for kinetic magnetism [74], where magnetic order arises from the path interference of a hole or doublon [75]. (2) What is the nature of the ground state? Ising spins exhibit glassy behavior on small-world and scale-free graphs [8, 76, 77, 78]. Can quantum fluctuations in the Heisenberg model stabilize a spin liquid in the absence of lattice symmetry [79, 80]? Answering this question would require a more detailed examination of the spin correlations. While we generally find antiferromagnetic two-site correlations, the correlation length grows with disassortativity, and second or third neighbors can be more strongly correlated than adjacent spins on hubbed networks (see End Matter). (3) How are the dynamical and ergodicity properties affected by network topology [23, 24]? (4) Is the magnetic order sensitive to community structures or other motifs [81] that lead to hierarchical or explosive synchronization in classical networks [4, 13, 11, 12, 82]?
Acknowledgements.
Acknowledgments—We thank Sitabhra Sinha, Sthitadhi Roy, Claudio Castelnovo, and Subhro Bhattacharjee for useful discussions and comments. We thank the National Supercomputing Mission (NSM) for providing computing resources of “PARAM Utkarsh” at CDAC Knowledge Park Bangalore, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. Data availability—The data and source codes that support our findings are openly available [83, 84].
@@ Acknowledgements.
Preethi Gopalakrishnan and Shovan Dutta
Appendix I Random trees
A tree has a bipartite structure [Fig. S1(a)], for which $S_{\text{total}}=|N_{A}-N_{B}|/2$, where $N_{A}$ and $N_{B}$ are the numbers of spins in the two sublattices [1]. Figure S1(b) shows that both the average and the spread of $S_{\text{total}}$ over random trees with $N$ sites and N1N-1 bonds scale as $\sqrt{N}$.
Figure S1: (a) Tree with $N=30$ sites. Horizontal lines show the two sublattices. (b) The average and the standard deviation of $S_{\text{total}}$ over ensembles of $5000$ random trees. Dashed lines show the fits $0.23\sqrt{N}$ and $0.17\sqrt{N}$, respectively.
Appendix II Random regular graphs
In the main text we showed that a random regular graph, where each node has $k=4$ neighbors, almost always has $S_{\text{total}}=0$ in ground state. Figure S2 shows the same holds for $k=3$ and $k=5$, further demonstrating that heterogeneity is essential for nonzero total spin.
@@ Appendix II Random regular graphs
Appendix III Quantum vs. classical $S_{\text{total}}$
Figure S3: (Top panel) Correlation and (bottom panel) root-mean-square deviation between the net magnetizations of classical and quantum spins for random graphs (left column) and copy-model graphs (right column) with $N=30$ sites.
Appendix IV Exact solution for the wheel graph
@@
In order to lower energy $\mathbf{S}_{b}$ and $\mathbf{S}_{c}$ point in opposite directions. Without loss of generality we take $\mathbf{S}_{b}$ along +z+z and $\mathbf{S}_{c}$ along z-z. The maximum magnitude of $\mathbf{S}_{c}$ is $N_{c}S$, where $S$ is the magnitude of each spin. To compare with qubits we set $S=1/2$ and write $\mathbf{S}_{c}=-(N_{c}/2)u\hat{\mathbf{z}}$, where $u\in[0,1]$. This central spin acts as an effective magnetic field for the Heisenberg ring [see Eq. (S1)], whose thermodynamic properties can be solved exactly [2]. The field wants to polarize the spins along +z+z, whereas the bonds prefer a Néel-type antiferromagnet. In the ground state, the spins orient along the polar and azimuthal angles
@@
where $\theta$ goes from 0 to $\pi/2$ as the field decreases, $\phi_{0}$ is an arbitrary constant, and $\alpha=\big[1-(-1)^{N_{b}}\big]\pi/(2N_{b})$. This canted antiferromagnet has energy
@@
$u=v=0\quad$ $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}>1\;,$ (S4a)
$u=v=1\quad$ $\text{if }\tilde{J}_{c}<1\text{ and }\tilde{J}_{b}^{\prime}<1\;,$ (S4b)
$u=1,v=1/\tilde{J}_{b}^{\prime}\quad$ $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}<1\text{ and }\tilde{J}_{b}^{\prime}>1\;,$ (S4c)
$v=1,u=1/\tilde{J}_{c}\quad$ $\text{if }\tilde{J}_{c}\tilde{J}_{b}^{\prime}<1\text{ and }\tilde{J}_{c}>1\;,$ (S4d)
@@
Figures S4(b) and S4(c) show that the resulting phase diagram differs qualitatively for $N_{c}=2$ and $N_{c}=3$. This is because for odd $N_{c}$ one cannot have $S_{c}=0$, which leads to a stepwise fall of $S_{b}$ as $J_{b}$ is increased relative to $J$. To understand this feature more precisely we consider the energy as a function of $S_{b}$ for a given $S_{c}<S_{b}$,
@@
which is seen in Fig. S4(b) for $S_{c}=1$ and in Fig. S4(c) for $S_{c}=3/2,1/2$. No such divergence is present for classical spins [Fig. S4(a)]. This disparity also arises for a Heisenberg ring in a classical magnetic field [6, 3, 7, 2].
[…]